A fence is 692m long. It is cut into three parts, two parts are the same length, and the third is 26m longer than either of the other two. Find the length of each part.
Answer by addingup(3677)

Let the two short parts be 2x. And the long part is x+26:
2x + x + 26 = 692 Add 2x + x on left and subtract 26, both sides
3x = 666 divide both sides by 3
x = 222
Each of the short parts is 222 long and the long part is 222 + 26 = 248m
Proof: 222 + 222 + 248 = 692